OpenStudy (golferbridget1999):

A circle has its center at (−1, 2) and a radius of 3 units. What is the equation of the circle? (x − 1)2 + (y + 2)2 = 3 (x + 1)2 + (y − 2)2 = 3 (x + 1)2 + (y + 2)2 = 9 (x + 1)2 + (y − 2)2 = 9

OpenStudy (amorfide):

standard equation of a circle \[(x+a)^{2}+(y+b)^{2}=r^2\] if we let x+a=0 solve for x x=-a and y+b=0 solve for y y=-b this means our centre is (-a, -b) so if you had a centre (2,3) and a radius 5 then x=2, so x-2=0 y=3, so y-3=0 you would substitute into the standard equation of a circle to get \[(x-2)^{2}+(y-3)^{2}=5^{2}\] so can you use these same steps to solve your question?